# To Trust or Not to Trust: On Calibration in ML-based Resource Allocation for Wireless Networks

### v1

@@ III-A Main Results
$\hskip 1.00006pt+\mathbb{E}[Q^{c}\!\mid Q^{c}\!\leq q_{th}]\!\left(\!1-\!(1-\!F(q_{th}))^{|\mathcal{R}|-1}\!\right)\!.$ (18)
@@
This corollary further simplifies the process of selecting an appropriate $q_{th}$. For example, (19) tells us that if the model is well-calibrated, all one needs to do is select a $q_{th}$ equal to their desired OP, and provided there are sufficient resources, they should achieve a system OP that is better than their desired OP.
Note that although calibration improves the alignment between predicted confidence and actual accuracy, it does not increase the informativeness of the predictor, as it does not introduce new information. As a result, post-processing calibration cannot improve the system’s minimum achievable OP, as formalized in the following theorem.
@@
$\min_{q_{th}}P^{c}_{|\mathcal{R}|}(\gamma_{th},q_{th})\geq\min_{q_{th}}P_{|\mathcal{R}|}(\gamma_{th},q_{th})$ (21)
@@
We now come to the final theoretical result of our work which introduces the concept of monotonicity that the accuracy-confidence function must satisfy for the predictor to be useful (i.e., satisfying (15)).
@@
Given a threshold level $q_{th}$, a predictor is useful (i.e., it satisfies (15)) if the inequality $\mathbb{E}\left[A(Q^{\star})|Q^{\star}\leq q_{th}\right]$ $<$ $P$ is satisfied. Furthermore, $\mathbb{E}\left[A(Q^{\star})|Q^{\star}\leq q_{th}\right]<P$ is satisfied if and only if
@@
See Appendix A-C
This theorem shows us that the conditional expectation of the predictor’s accuracy-confidence function should be smaller when conditioned on values less than or equal to $q_{th}$ than when conditioned on values greater than $q_{th}$. A trivial class of accuracy-confidence functions that achieve this are those that are monotonically increasing. As an example, if $A(q)=q$ for all values of $q$, as is the case for perfectly calibrated predictors, then $A$ is a monotonically increasing function and so the corresponding predictor necessarily improves the OP. There may be examples of strongly over-confident and strongly under-confident predictors that also satisfy this monotonicity property (see Fig.1 (top)). Importantly, $A(q)$ being monotonically increasing is a sufficient but not necessary condition required for a predictor to be useful. For example, it is possible for $A(q)$ to be non-monotonic in the regions $[0,q_{th})$ and $[q_{th},1]$ whilst also satisfying $A(q)>A(q^{\prime})$ for $q^{\prime}\in[0,q_{th})$ and $q\in[q_{th},1]$. Such a function would also satisfy (22) (see Fig.2 (a)). Conversely, we can also see that if a predictor’s accuracy-confidence function is monotonically decreasing, then it will necessarily hinder the system’s OP (see Fig.2 (b)). Again, using a similar example as above, this is a sufficient but not necessary property of a predictor that hinders performance.
@@ IV Methodology
This section describes the data-generation process and outlines the key metrics used to evaluate calibration performance of the outage predictor. It also provides an overview of the post-processing calibration techniques implemented in this study, along with relevant implementation details.
@@ IV-A Data Generation
The training/testing dataset for the channel is synthetically obtained as follows. We start by generating a zero-mean complex Gaussian vector given by: $G[0]=[g_{1},g_{2},\dots,g_{v}]$ at $t=0$, with $v=1024$, where each element has a variance of $\frac{1}{v}$. To simulate the motion of a mobile user, at each time step $t$, we apply a per-element, independent and uniformly distributed phase shifts $\theta_{it}\sim\text{Unif}\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$. At time $t=1$, we have:
@@
where $\mathrm{J_{0}(\cdot)}$ is the zeroth-order Bessel function of the first kind, and (27) follows from the characteristic function of the uniform distribution. This model is in-line with the widely accepted Clarke’s 2D model [^2] Additionally, channel data generated using Clarke’s 3D model [75] was used, and the results discussed in the subsequent subsections apply equally to this dataset. [76].
TABLE I: DQN-LSTM Model Architecture and Hyperparameters
Architecture Specification
Layer 1: LSTM layer 32 hidden units
Layer 2: Dense layer 10 units with PReLU activation
Layer 3: Dense layer 1 unit with sigmoid activation
Hyperparameter Values
Hidden units 32
Epochs 30
Epoch size 150
Input sequence length ($k$) 100
Output sequence length ($l$) 10
Output dimension 1
Learning rate 0.001
Discount factor 0.9
Epsilon Decaying strategy
Rewards
1 (correct prediction);
-1 (otherwise)
@@ IV-B Training and Evaluation
We utilize a Deep Q-Network (DQN) incorporating an LSTM layer, following the implementation in [77], and train it on the channel data using both the OLF from [46] ($\alpha=10$) and the BCE loss function. The predictor is developed using TensorFlow’s Keras API and trained with the ADAM optimizer. Table I details the architecture and hyperparameters of the DQN-LSTM predictor employed in our experiments. We train our predictor on a dataset comprising 4,500 × (number of resources) samples. For a 4-resource system, this corresponds to 18,000 training samples, with the dataset size scaling proportionally (e.g., 27,000 samples for a 6-resource system). The training and validation datasets are identical in size, and testing is performed on 11,000 instances of the $|\mathcal{R}|$-resource system to achieve statistically robust results. The predictor is retrained 10 times, and the average performance is used as a single data point for analysis. Each experiment generates sequences of $k+l$ channel samples per resource, where $k=100$, $l=10$, and $\xi=0.1$ radians. The first $k$ samples are provided as input to the DQN-LSTM model, while the last $l$ samples, $\accentset{\rightarrow}{H}_{j}^{l}{(t)}$, are used to determine whether communication can be supported at a specified rate. On a desktop computer equipped with a 12th Gen Intel(R) Core(TM) i9-12900HK, 2500 MHz, 14 Core(s) and 20 Logical processor(s), the training of a single DQN-LSTM model takes approximately 5 minutes.
The reliability diagrams presented in Section V are constructed using logarithmic binning as defined in [2, Eq. (7)]:
@@
where $b=\sqrt{10}$. These reliability plots represent the average calibration performance of our outage predictor over 10 runs, with each run evaluating calibration on 10,000 samples.
To empirically demonstrate the key properties of the OP established under perfect calibration in Section III, we apply two widely used post-processing calibration techniques: Platt scaling and isotonic regression. Their effectiveness is evaluated using standard calibration metrics. Specifically, we use Negative log-likelihood (NLL) [1] and Brier score (BS) [78, 79], which assess how well predicted probabilities align with actual outcomes for individual predictions. NLL, also known as cross-entropy loss [80], measures the alignment of predicted probabilities with true labels [66]. The BS captures the mean squared difference between predicted probabilities and true outcomes, with values $\in[0,1]$. For both the metrics, lower values indicate better calibration.
@@
    where $z_{i}$ is the logit, $m$ and $n$ are parameters learned during calibration, and $\sigma(\cdot)$ is the sigmoid function. Since the DQN-LSTM outputs probabilities through a sigmoid activation, we first convert them to logits using the logit function:
@@ V Results and Demonstration
This section presents numerical results that demonstrate how the output of a calibrated predictor relates to the system’s OP and helps identify an appropriate $q_{th}$ to meet reliability targets. We begin by an evaluation of the predictor’s calibration using post-processing methods and reliability diagrams with logarithmic binning. This is followed by a robust simulation-based analysis that demonstrate the theoretical results established in Theorems 1 and 2. Finally, we present a proof of concept for the resource allocation technique introduced in Section II.
Figure 3: Histogram-based reliability diagrams of our predictor utilizing OLF and BCE for a 4-resource system with a rate threshold $\gamma_{th}$ of 0.5. Specifically, plots are presented for $\mathtt{SNR}$ values of (a) 2 dB and (b) 8 dB for predictors trained with OLF and (c) 2 dB and (d) 8 dB for predictors trained with BCE with various calibration methods.
@@ V-A Results
Fig. 3 presents the histogram-based reliability plots using logarithmic binning for our outage predictor with a 4-resource system where $\gamma_{th}=0.5$ and $q_{th}=0.5$. Fig. 3 (a) and (b) show logarithmic reliability plots for predictors trained with OLF at an $\mathtt{SNR}$ of 2 dB and 8 dB, respectively, while Fig. 3 (c) and (d) present the corresponding logarithmic reliability plots for predictors trained with BCE. Furthermore, Table II presents the NLL and BS values for predictors trained with OLF and BCE, comparing the performance of calibrated and uncalibrated predictors across different $\mathtt{SNR}$ levels.
TABLE II: NLL and Brier Score (BS) Values for different $q_{th}$ and $\mathtt{SNR}$ Levels: Uncalibrated vs. Calibrated
SNR OLF BCE
@@
Figure 4: Comparison of $\!P^{\star}_{\infty}(\gamma_{th},q_{th})$ and $\mathbb{E}[Q^{\star}\!\mid Q^{\star}\leq q_{th}]$ across different $q_{th}$ values, in a 4-resource system using OLF with $\gamma_{th}\!=2.0$ and $\mathtt{SNR}=6\mathrm{dB}$.
Figure 5: Comparison of $P$ and $\mathbb{E}[Q^{\star}]$ across different numbers of resources $\mathcal{R}$, using OLF with $\gamma_{th}=1.0$ and $\mathtt{SNR}=6~{}\mathrm{dB}$.
From Fig. 3, it is evident that the application of both the calibration methods significantly enhances the calibration performance of the outage predictor. This improvement is not only observed visually but also supported by the lower NLL and BS values achieved post calibration. To better understand their impact, we assess the effectiveness of each calibration method on the OLF-trained outage predictor below.
@@
    Platt Scaling: This method improves calibration, as shown in Fig. 3 (a) for $q_{th}=0.5$ and $\mathtt{SNR}=2~{}\mathrm{dB}$, where the grey bars represent the uncalibrated results and the dark blue bars represent the Platt scaled results. The calibration improvement is also reflected in the reduced NLL and BS values after calibration (see Table II), with NLL decreasing from 0.33 (uncalibrated) to 0.12 (Platt scaled) and BS dropping from 0.09 (uncalibrated) to 0.05 (Platt scaled). In general, this method demonstrates its effectiveness primarily at higher confidence levels, particularly within the range of $10^{-2}$ to $1$.
@@
    Isotonic Regression: This method provides significant calibration improvements through non-parametric adjustments to predicted probabilities. For example, in Fig. 3 (b), for $q_{th}=0.5$ and $\mathtt{SNR}=8~{}\mathrm{dB}$, the dark red bars align closely with the perfect calibration line, illustrating its effectiveness. This is further validated by a notable reduction in NLL from 0.18 (uncalibrated) to 0.03 (isotonic scaled) and in BS from 0.06 (uncalibrated) to 0.01 (isotonic scaled). It can be observed that this method proves to be effective even at lower confidence levels.
@@
Figure 6: $P^{\star}_{|\mathcal{R}|}(\gamma_{th},q_{th})$ vs $q_{th}$ for a 4-resource system using OLF with (a) $\gamma_{th}~{}=0.5$, at $\mathtt{SNR}$ of 2 dB and (b) $\gamma_{th}=1.0$ at $\mathtt{SNR}$ of 8 dB.
@@
Figure 7: $P^{\star}_{|\mathcal{R}|}(\gamma_{th},q_{th})$ vs $q_{th}$ for a 4-resource system using BCE with (a) $\gamma_{th}~{}=0.5$, at $\mathtt{SNR}$ of 2 dB and (b) $\gamma_{th}=1.0$ at $\mathtt{SNR}$ of 8 dB.
@@
From Fig. 3 (c), and (d), it can be seen that the predictors trained with BCE display better calibration in their uncalibrated state compared to those trained with OLF. This is because BCE-based training is independent of $q_{th}$, enabling it to model the overall channel distribution. Post calibration, the predictors trained with BCE showed minimal changes, as reflected by the NLL values and BS (see Table II).
@@
Figs. 4 and 5 illustrate the key results [^3] These plots do not compare the outage performance of OLF with BCE as $q_{th}$ is not a hyperparameter for BCE-based training. discussed in Section III, demonstrating the behaviour established for a perfectly calibrated predictor in Theorem 1. For this analysis, we apply isotonic regression, as it consistently achieves better calibration performance, according to the NLL and BS metrics, compared to Platt scaling. In particular, Fig. 4 shows the OP for an infinite resource limit, $P^{\star}_{\infty}(\gamma_{th},q_{th})$ alongside the expected confidence of the predictor when its output is below the threshold $q_{th}$, i.e., $\mathbb{E}[Q^{\star}\!\mid\!Q^{\star}\leq q_{th}]$. The figure compares calibrated and uncalibrated predictors over a range of $q_{th}$ values, using a system with 4 resources, $\gamma_{th}=2.0$, and $\mathtt{SNR}=6~{}\mathrm{dB}$. It can be observed that these curves align closely for the calibrated predictor when isotonic regression is applied, consistent with the behavior described in Theorem 1, Eq. (17).
Furthermore, this result helps in selecting an appropriate $q_{th}$ to meet a target system OP. For instance, a service provider might have a service level agreement (SLA) requiring the OP to be no greater than $0.01$ (i.e., a user should experience an outage no more than once every 100 transmissions). To satisfy this requirement, the provider can consult the calibrated outage predictor [^4] The NN used here is lightweight, enabling fast training, in contrast to large-scale models with extensive parameter counts.
This example applies only to a calibrated predictor, for which (17) holds. In the uncalibrated case, the output confidence does not reflect the true OP, this equation no longer holds, and $q_{th}$ must be chosen via empirical search.
@@
. Given $\gamma_{th}=2.0$, $\mathtt{SNR}=6~{}\mathrm{dB}$, and a sufficient number of resources, Fig. 4 shows that setting $q_{th}=0.2$ with a well-calibrated predictor satisfies this SLA. Additionally, we observe that the calibrated and uncalibrated OLF-trained predictors both achieve the same minimum OP but at different $q_{th}$ values. The shift in the $q_{th}$ value where the calibrated predictor achieves its minimum OP is a consequence of calibration, which adjusts the predicted confidence scores to better reflect true OPs thereby rescaling the threshold at which the minimum OP occurs.

### v3

@@ III-A Main Results
$\thinspace+\mathbb{E}[Q^{c}\!\mid Q^{c}\!\leq q_{th}]\!\left(\!1-\!(1-\!F(q_{th}))^{|\mathcal{R}|-1}\!\right).$ (19)
@@
This corollary further simplifies the process of selecting an appropriate $q_{th}$. Equation (20) tells us that if the model is well-calibrated, all one needs to do is select a $q_{th}$ equal to their desired OP, and provided there are sufficient resources, they should achieve a system OP that is better than their desired OP. For example, if a system requires an OP of 0.01, setting $q_{th}=0.01$ ensures that the system OP will be better than the desired OP.
Note that although calibration improves the alignment between predicted confidence and actual accuracy, it does not increase the informativeness of the predictor. Post-processing calibration methods adjust the output scores using a held-out validation set, but they do not alter the underlying model or incorporate any new information. As a result, such methods cannot reduce the system’s minimum achievable OP, in accordance with the data processing inequality, as formalized in the following theorem.
@@
$\min_{q_{th}}P^{c}_{|\mathcal{R}|}(\gamma_{th},q_{th})\geq\min_{q_{th}}P_{|\mathcal{R}|}(\gamma_{th},q_{th}).$ (22)
@@
We now come to the final theoretical result of our work, which introduces the concept of monotonicity that the accuracy-confidence function must satisfy for the predictor to be useful (i.e., satisfying (16)).
@@
Given a threshold level $q_{th}$, a predictor is useful (i.e., it satisfies (16)) if the inequality $\mathbb{E}\left[A(Q^{\star})|Q^{\star}\leq q_{th}\right]$ << $P$ is satisfied. Furthermore, $\mathbb{E}\left[A(Q^{\star})|Q^{\star}\leq q_{th}\right]<P$ is satisfied if and only if
@@
See Appendix A-C. ∎
This theorem shows us that the conditional expectation of the predictor’s accuracy-confidence function should be smaller when conditioned on values less than or equal to $q_{th}$ than when conditioned on values greater than $q_{th}$. A trivial class of accuracy-confidence functions that achieve this are those that are monotonically increasing. As an example, if $A(q)=q$ for all values of $q$, as is the case for perfectly calibrated predictors, then $A$ is a monotonically increasing function and so the corresponding predictor necessarily improves the OP. There may be examples of strongly over-confident and strongly under-confident predictors that also satisfy this monotonicity property (see Fig. 1 (top)). Importantly, $A(q)$ being monotonically increasing is a sufficient but not necessary condition required for a predictor to be useful. For example, it is possible for $A(q)$ to be non-monotonic in the regions $0,q_{th})$ and $q_{th},1]$ whilst also satisfying $A(q)>A(q^{\prime})$ for $q^{\prime}\in[0,q_{th})$ and $q\in[q_{th},1]$. Such a function would also satisfy (23) (see Fig. 2 (a)). Conversely, we can also see that if a predictor’s accuracy-confidence function is monotonically decreasing, then it will necessarily hinder the system’s OP (see Fig. 2 (b)). Again, using a similar example as above, this is a sufficient but not necessary property of a predictor that hinders performance.
@@ IV Methodology
This section describes the data-generation process and outlines the key metrics used to evaluate the calibration performance of the outage predictor. It also provides an overview of the post-processing calibration techniques implemented in this study, along with relevant implementation details.
TABLE I: Model Architectures and Hyperparameters
Component Specification
Architecture (common to both models)
Layer 1: LSTM layer 1 layer with 32 hidden units
Layer 2: Dense layer 10 units with PReLU activation
Layer 3: Dense layer 1 unit with sigmoid activation
Training Hyperparameters
Learning rate 0.001
Number of epochs 30
Epoch size 150
Batch size Number of resources
Input sequence length ($k$) 100
Output sequence length ($l$) 10
Output dimension 1
DQN-specific Parameters
Discount factor 0.9
Epsilon Decaying strategy
Rewards 1 (correct prediction); -1 (otherwise)
OLF Parameters
$\alpha$ 10
$q_{th}$ 0.5
Figure 3: System-level overview of the ML-based outage prediction framework. The predictor takes dynamic wireless channel data for resource $j\in\mathcal{R}$ as input and produces uncalibrated confidence estimates. Post-processing calibration yields calibrated confidence values, which are then used for threshold-based resource selection.
@@ IV-A Data Generation
The training/testing dataset for the channel is synthetically obtained as follows. We start by generating a zero-mean complex Gaussian vector given by: $G[0]=[g_{1},g_{2},\dots,g_{v}]$ at $t=0$, with $v=1024$, where each element has a variance of $\frac{1}{v}$. To simulate the motion of a mobile user, at each time step $t$, we apply a per-element, independent and uniformly distributed phase shift $\theta_{it}\sim\text{Unif}\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$. At time $t=1$, we have:
@@
where $\mathrm{J_{0}(\cdot)}$ is the zeroth-order Bessel function of the first kind, and (28) follows from the characteristic function of the uniform distribution. This model is in line with the widely accepted Clarke’s 2D model [^2] Additionally, channel data was generated using Clarke’s 3D model [76], where the autocorrelation is given by $\mathrm{sinc}(\cdot)$, in contrast to Clarke’s 2D model used in this paper, where the autocorrelation is given by $\mathrm{J_{0}}(\cdot)$. This was done to support the general validity of our theoretical results across different input data distributions. The findings discussed in the following subsections were found to be consistent and equally applicable to this dataset. [77], where $\xi$ represents the displacement of the mobile receiver, thereby capturing the temporal correlation introduced by user mobility. This model ensures that the simulated fading process reflects realistic channel dynamics, enabling a meaningful evaluation of the outage predictor’s calibration performance.
@@ IV-B Training and Evaluation
We utilize an ML-based outage predictor and compare two model architectures: a Deep Q-Network (DQN) incorporating an LSTM layer, as implemented in [78], and a standalone LSTM model, as implemented in [48]. Both predictors are trained on channel data generated using Clarke’s 2D model, employing the OLF from [48] as well as the BCE loss function. The predictors are developed using TensorFlow’s Keras API and trained with the ADAM optimizer. Table I details the architecture and hyperparameters of both the ML models used in our experiments.
We train our predictors on a dataset comprising 4,500 × (number of resources) samples. For a 4-resource system, this corresponds to 18,000 training samples, with the dataset size scaling proportionally (e.g., 27,000 samples for a 6-resource system). The training and validation datasets are identical in size, and testing is performed on 11,000 instances of the $|\mathcal{R}|$-resource system to achieve statistically robust results. The predictor is retrained 10 times, and the average performance is used as a single data point for analysis. Each experiment generates sequences of $k+l$ channel samples per resource, where $k=100$, $l=10$, and $\xi=0.1$ radians. The first $k$ samples are provided as input to the ML model, while the last $l$ samples, $\accentset{\rightarrow}{H}_{j}^{l}{(t)}$, are used to determine whether communication can be supported at a specified rate. On a desktop computer equipped with a 12th Gen Intel(R) Core(TM) i9-12900HK, 2500 MHz, 14 Core(s) and 20 Logical processor(s), the training of a single predictor takes approximately 5 minutes.
While post-processing calibration helps improve the reliability of predicted confidence scores, it introduces a slight computational overhead since it is applied separately after model training. However, the NN used in this study is lightweight and enables fast training, so the additional cost of calibration remains minimal, especially when compared to large-scale models with extensive parameter counts. To reduce computational overhead in practical implementations, calibration objectives can be integrated directly into the training process [79, 80]. By eliminating the need for separate post-processing calibration, this approach can be more suitable for deployment in energy-constrained or real-time systems.
The reliability diagrams presented in Section V are constructed using logarithmic binning as defined in [4, Eq. (7)]:
@@
where $b=\sqrt{10}$. This binning scheme enables the assessment of calibration behavior across several orders of magnitude in predicted confidence, with enhanced resolution in the low-confidence regions. This feature is particularly relevant in wireless communications, where systems often operate in the very low regions of performance metrics such as OP, bit/symbol error rates, etc. These regimes correspond to high-reliability operation, where fine-grained calibration assessment becomes crucial. These reliability plots represent the average calibration performance of our outage predictor over 10 runs, with each run evaluating calibration on 10,000 samples.
To empirically demonstrate the key properties of the OP established under perfect calibration in Section III, we apply two widely used post-processing calibration techniques: Platt scaling and isotonic regression. Their effectiveness in improving calibration is evaluated using two standard metrics: Negative Log-Likelihood (NLL) [3] and Brier Score (BS) [81, 82]. NLL, also known as cross-entropy loss [83], measures the alignment of predicted probabilities with true labels [68]. The BS captures the mean squared difference between predicted probabilities and true outcomes, with values $\in[0,1]$. For both the metrics, lower values indicate better calibration.
@@
    where $z_{i}$ is the logit, $m$ and $n$ are parameters learned during calibration, and $\sigma(\cdot)$ is the sigmoid function. Since the ML models used here outputs probabilities through a sigmoid activation, we first convert them to logits using the logit function:
@@
Figure 4: Flowchart illustrating the key steps required to reproduce the proposed calibration-aware ML-based resource allocation framework.
Fig. 3 shows a system-level overview of the ML-based outage prediction framework. The predictor processes dynamic wireless channel data for resource $j\in\mathcal{R}$ and outputs uncalibrated confidence estimates. These estimates are then calibrated through post-processing calibration methods, resulting in confidence values that are used to select resources based on a predefined classification threshold. Furthermore, a flowchart outlining the key steps involved in implementing the proposed ML-based resource allocation framework, including post-processing calibration, is shown in Fig. 4.
Our study employs offline, post-processing calibration methods such as Platt scaling and isotonic regression. However, these static approaches may deteriorate under distribution shifts [84]. These shifts can be mitigated using online adaptation, which updates NN parameters during deployment to adapt to evolving data distributions, or through robust strategies that improve calibration performance across diverse conditions [85]. Examples include architectural designs such as Annealing Double-Head networks [86], gradient-based updates via differentiable or approximated solvers, and gradient-free schemes like ensemble Kalman inversion [87], all of which help sustain calibration under shifting conditions.
@@ V Results and Demonstration
This section presents numerical results that demonstrate how the output of a calibrated predictor relates to the system’s OP and helps identify an appropriate $q_{th}$ to meet reliability targets. We begin by evaluating the predictor’s calibration using post-processing methods and reliability diagrams with logarithmic binning. This is followed by a robust simulation-based analysis that demonstrates the theoretical results established in Theorems 1 and 2. Finally, we present a proof of concept for the resource allocation technique introduced in Section II.
Figure 5: Histogram-based reliability diagrams of our predictor utilizing OLF and BCE for a 4-resource system with a rate threshold $\gamma_{th}$ of 0.5, using both LSTM and DQN-LSTM model architectures. Specifically, plots are presented for $\mathtt{SNR}$ values of (a) 2 dB and (b) 8 dB, for predictors trained with OLF, and (c) 2 dB and (d) 8 dB, for predictors trained with BCE, each with various calibration methods applied.
@@ V-A Results
Fig. 5 presents the histogram-based reliability plots using logarithmic binning for our outage predictor with a 4-resource system where $\gamma_{th}=0.5$ and $q_{th}=0.5$. The results are shown for both the LSTM and DQN-LSTM model architectures. Fig. 5 (a) and (b) show logarithmic reliability plots for predictors trained with OLF at an $\mathtt{SNR}$ of 2 dB and 8 dB, respectively, while Fig. 5 (c) and (d) present the corresponding logarithmic reliability plots for predictors trained with BCE. Furthermore, Table II presents the NLL and BS values for predictors trained with OLF and BCE, comparing the performance of calibrated and uncalibrated predictors across different $\mathtt{SNR}$ levels for both model architectures.
TABLE II: NLL and BS Values across ML models and loss functions at different $\mathtt{SNR}$ Levels: Uncalibrated vs. Calibrated
SNR OLF (DQN-LSTM) BCE (DQN-LSTM)
@@
SNR OLF (LSTM) BCE (LSTM)
NLL BS NLL BS
2 dB 0.36 0.10 0.08 0.09 0.03 0.02 0.08 0.08 0.07 0.03 0.03 0.03
8 dB 0.22 0.05 0.02 0.05 0.01 0.01 0.04 0.03 0.03 0.03 0.03 0.03
From Fig. 5, it is evident that the application of both the calibration methods significantly enhances the calibration performance of the outage predictor. This improvement is not only observed visually but also supported by the lower NLL and BS values achieved after calibration. Notably, both the LSTM and DQN-LSTM model architectures achieve similar performance, highlighting the effectiveness of the calibration methods across both models. To better understand their impact, we assess the effectiveness of each calibration method on the OLF-trained outage predictor below.
@@
    Platt Scaling: This method improves calibration, as shown in Fig. 5 (a) for $\gamma_{th}=0.5$ and $\mathtt{SNR}=2~\mathrm{dB}$, where the dark grey bars represent the uncalibrated results and the dark blue bars represent the Platt scaled results, obtained using the DQN-LSTM model. The calibration improvement is also reflected in the reduced NLL and BS values after calibration (see Table II), with NLL decreasing from 0.33 (uncalibrated) to 0.12 (Platt scaled) and BS dropping from 0.09 (uncalibrated) to 0.05 (Platt scaled). In general, this method demonstrates its effectiveness primarily at higher confidence levels, particularly within the range of $10^{-2}$ to $1$.
@@
    Isotonic Regression: This method provides significant calibration improvements through non-parametric adjustments to predicted probabilities. For example, in Fig. 5 (b), for $\gamma_{th}=0.5$ and $\mathtt{SNR}=8~\mathrm{dB}$, the light grey bars represent the uncalibrated results and the light red bars represent the isotonic scaled results, using the LSTM model. Here, the light red bars align closely with the perfect calibration line, illustrating its effectiveness. This is further validated by a notable reduction in NLL from 0.22 (uncalibrated) to 0.02 (isotonic scaled) and in BS from 0.05 (uncalibrated) to 0.01 (isotonic scaled). It can be observed that this method proves to be effective even at lower confidence levels.
@@
Figure 6: $P^{\star}_{|\mathcal{R}|}(\gamma_{th},q_{th})$ vs $q_{th}$ for a 4-resource system using OLF, evaluated for both LSTM and DQN-LSTM architectures, with (a) $\gamma_{th}~=0.5$, at $\mathtt{SNR}$ of 2 dB and (b) $\gamma_{th}=1.0$, at $\mathtt{SNR}$ of 8 dB.
@@
Figure 7: $P^{\star}_{|\mathcal{R}|}(\gamma_{th},q_{th})$ vs $q_{th}$ for a 4-resource system using BCE, evaluated for both LSTM and DQN-LSTM architectures, with (a) $\gamma_{th}~=0.5$, at $\mathtt{SNR}$ of 2 dB and (b) $\gamma_{th}=1.0$, at $\mathtt{SNR}$ of 8 dB.
Platt scaling, being parametric, is efficient and less prone to overfitting when the held-out validation data is limited, whereas isotonic regression, as a non-parametric method, offers greater flexibility and achieves better calibration with sufficient data. In our setting, where accuracy across a broader confidence range is critical, isotonic regression consistently delivers superior results, as reflected by its lower NLL and BS values.
@@
From Fig. 5 (c), and (d), it can be seen that the predictors trained with BCE display better calibration in their uncalibrated state compared to those trained with OLF. This is because BCE-based training is independent of $q_{th}$, enabling it to model the overall channel distribution. Post calibration, the predictors trained with BCE showed minimal changes, as reflected by the NLL and BS values (see Table II).
Figs. 6 and 7 presents the OP for a 4-resource system using OLF and BCE, respectively, across a range of $q_{th}$ values, for (a) $\gamma_{th}=0.5$ at $\mathtt{SNR}=2~\mathrm{dB}$ and (b) $\gamma_{th}=1.0$ at $\mathtt{SNR}~=~8~\mathrm{dB}$, using both LSTM and DQN-LSTM model architectures. Unlike Fig. 5 from [48], these plots use $q_{th}$ values on a logarithmic scale and also present the OP obtained using the calibrated predictors. It can be seen that the OP obtained using OLF and BCE exhibit a ‘U’-shaped curve, indicating that both very small and very large $q_{th}$ values are not good. This is because very small $q_{th}$ values result in the user almost always being allocated the final resource, while very large $q_{th}$ values lead to allocation to the first resource. Notably, at $q_{th}\!=\!0.5$, the uncalibrated OLF-trained predictor outperforms the BCE-trained ones. As $q_{th}$ decreases, this performance gap narrows, and in some cases, BCE slightly outperforms OLF at lower $q_{th}$.
It can also be seen from Fig. 6 that the calibrated and uncalibrated OLF-trained predictors both achieve the same minimum OP but at different $q_{th}$ values. For instance, in a 4-resource system with $\gamma_{th}=0.5$ and $\mathtt{SNR}=2~\mathrm{dB}$ (see Fig. 6 (a)), the uncalibrated predictor trained with OLF achieves its minimum OP at $q_{th}=0.5$, which matches its training threshold, while the calibrated predictor reaches the same minimum OP at $q_{th}=10^{-3}$. This observation aligns with the result of Theorem 2, which establishes that post-processing calibration cannot reduce the minimum achievable OP. It is important to note that this observation is not reflected in Fig. 7, because BCE-trained predictors are inherently well calibrated, as evidenced by Fig. 5 (c) and (d), and Table II. As previously mentioned, the distinction between the two training approaches highlights key differences in their calibration behavior. BCE-trained predictors generally exhibit better inherent calibration, as their training is independent of $q_{th}$ and models the entire channel distribution, whereas OLF is specifically designed to optimize outage performance at a fixed classification threshold.
Figure 8: Comparison of $P^{\star}_{\infty}(\gamma_{th},q_{th})$ and $\mathbb{E}[Q^{\star}\!\mid Q^{\star}\leq q_{th}]$ across different $q_{th}$ values, in a 4-resource system using OLF, evaluated using both LSTM and DQN-LSTM architectures, with $\gamma_{th}\!=2.0$ and $\mathtt{SNR}=6\mathrm{dB}$.
Figure 9: Comparison of $P$ and $\mathbb{E}[Q^{\star}]$ across different numbers of resources $\mathcal{R}$, using OLF, evaluated using both LSTM and DQN-LSTM architectures, with $\gamma_{th}=1.0$ and $\mathtt{SNR}=6~\mathrm{dB}$.
@@
Figs. 8 and 9 illustrate the key results [^3] These plots do not compare the outage performance of OLF with BCE, as BCE-trained predictors are already well calibrated in their uncalibrated state (see Fig. 5 (c) and (d), and Table II). discussed in Section III, demonstrating the behaviour established for a perfectly calibrated predictor in Theorem 1, using both LSTM and DQN-LSTM models. For this analysis, we apply isotonic regression, as it consistently achieves better calibration performance, according to the NLL and BS metrics, compared to Platt scaling. In particular, Fig. 8 shows the OP for an infinite resource limit, $P^{\star}_{\infty}(\gamma_{th},q_{th})$ alongside the expected confidence of the predictor when its output is below the threshold $q_{th}$, i.e., $\mathbb{E}[Q^{\star}\!\mid\!Q^{\star}\leq q_{th}]$. The figure compares calibrated and uncalibrated predictors over a range of $q_{th}$ values, using a system with 4 resources, $\gamma_{th}=2.0$, and $\mathtt{SNR}=6~\mathrm{dB}$. It can be observed that these curves align closely for the calibrated predictor when isotonic regression is applied, consistent with the behavior described in Theorem 1, Eq. (18).
@@
Furthermore, this result helps in selecting an appropriate $q_{th}$ to meet a target system OP. For instance, a service provider might have a service level agreement (SLA) requiring the OP to be no greater than $0.01$ (i.e., a user should experience an outage no more than once every 100 transmissions). To satisfy this requirement, the provider can consult the calibrated outage predictor [^4] This example applies only to a calibrated predictor, for which (18) holds. In the uncalibrated case, the output confidence does not reflect the true OP; this equation no longer holds, and $q_{th}$ must be chosen via empirical search.. Given $\gamma_{th}=2.0$, $\mathtt{SNR}=6~\mathrm{dB}$, and a sufficient number of resources, Fig. 8 shows that setting $q_{th}=0.2$ for a well-calibrated predictor satisfies this SLA. Additionally, we observe that both the calibrated and uncalibrated OLF-trained predictors achieve the same minimum OP, but at different $q_{th}$ values. The shift in the $q_{th}$ value where the calibrated predictor achieves its minimum OP is a consequence of calibration, which adjusts the predicted confidence scores to better reflect true OPs thereby rescaling the threshold at which the minimum OP occurs.